# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ III-C2 Applications
Multi-modal sensing has empowered various kinds of applications. Via the utilization of audio and visual data, the performance of speech recognition and synthesis [67, 167], human action classification [168], multimedia event detection [169], emotion recognition [68], can be greatly enhanced compared with using unimodal information. By utilizing the modality translation techniques, many media description applications have been well developed, including image and video captioning [69, 170], visual question-answering [171], etc. By creating structured representations of multi-modal data, multimedia content indexing, and retrieval applications are enabled, including automatic shot-boundary detection [172] and video summarization [173]. Furthermore, multi-modal data can enhance the understanding, analysis, and decision-making processes in the domain of healthcare, fitness, and general wellness [174, 175]. Autonomous systems utilize various kinds of sensory data, including radar, stereoscopic visible-light cameras, light detection and ranging (LiDAR), infrared cameras, a global positioning system (GPS), and audio for enhancing the automation capability [176, 177].
@@ III-D2 Key Techniques
  • Data Collection: In mobile crowdsensing systems, data collection requires the engagement of massive user participation, where some non-professional users are included. The target of data collection techniques is acquiring reliable and trustworthy data via sensor selection. There are two frameworks. One is the domain-specific framework designed for a dedicated task such as intelligent transportation and health care [33]. As a result, the performance of the specific task can be enhanced. The main challenge of such a framework lies in the co-existence of many tasks or domains degrading the efficiency and reliability of the overall crowdsensing system. The other is a general-purpose data collection framework. As its name suggests, it is designed to support many tasks at the same time by maximizing a set of parameters or minimizing the costs [179].
  • Data Storage: The practical storage location in crowdsensing systems is task-oriented and is classified into centralized storage and distributed storage. The former stores data in a single location, typically a cloud-based database, which is applied when significant processing or data aggregation is needed, e.g., in urban monitoring and price comparison [33]. This approach requires high communication throughput for data transmission and violates the data privacy of users but can enjoy a low processing latency. Distributed storage is employed for delay-tolerant applications such as mapping air quality and fog computing, but has low communication cost and can keep users’ data privacy [180]. The storage format can be structured and unstructured [33]. In a structured format, the data has a specific identifier and is easy to access and self-explanatory, e.g., the information of individuals stored in terms of their names or identifications. In an unstructured format, there is no specific identifier, leading to difficulty in recognizing the data, e.g., the storage of videos, audio, and images.
  • Signal Processing: Signal processing in crowdsensing refers to the manipulation and analysis of the collected data. It involves three sequential processes, i.e., pre-processing, analytics, and post-processing [33]. The main target of pre-processing is filtering and denoising by removing irrelevant and redundant information, suppressing noise, aggregating data, and compressing data size to save storage volume [75]. Analytics aims to extract meaningful information from the collected data. Machine learning and data mining are common approaches for information inference, pattern recognition, or future trend prediction. For real-time tasks such as emergency situation management, traffic monitoring, and unmanned vehicles, high computational resources are needed for fast decision-making and response [33]. After analytics, post-processing can be conducted in an offline way for statistical or predictive analysis [181].
@@
    • Task assignment: Task assignment involves two aspects. One is task partitioning, which divides complex crowdsensing tasks into smaller and manageable sub-tasks, that can be completed by different devices [76]. For example, in [182], the overall target area is divided into multiple distinct sub-areas, with each being assigned to one device for sensing. The spatial and temporal correlation of different devices’ sensory data are leveraged. In [183], K-means algorithm is used to partition a crowdsensing task into several sub-ones, each of which is allocated to one mobile device according to their sensing abilities. The similarity of the sub-tasks is further considered in [184] during task assignment.
      The other task assignment technique is task allocation, which involves distributing tasks to users with the goal of reducing resource costs and enhancing task performance. This technique is categorized into two groups, i.e., centralized methods and distributed methods. In the former, a central coordinator assigns the tasks to users or devices, leading to an enhanced systematic performance but requiring global information (see, e.g., [182, 185]). In the latter, each device makes their own decisions during the sensing task, resulting in better data privacy preservation (see, e.g., [186]).
    • Task scheduling: Task scheduling in mobile crowdsensing refers to the way that mobile devices collect and contribute their sensory data. In proactive scheduling, mobile devices actively obtain sensory data without a pre-assigned task and can decide where and when to contribute their data. Proactive scheduling is commonly conducted for people-centric tasks (see, e.g., [187, 188]). In reactive scheduling, the sensing task and target are clearly pre-defined, which is adopted in many works. For example, the authors in [189] propose a minimum energy multi-sensor task scheduling algorithm that coordinates multiple tasks for the same user under the targets of enhancing sensory data quality and minimizing energy consumption. A first-in-first-out scheduling strategy and an arbitrary deadline (AD) scheduling strategy are respectively proposed for offline and online settings in [190]. A deep reinforcement-learning-based real-time scheduling algorithm is proposed in [191].
@@ III-D3 Incentive Mechanisms
The crucial condition for the success of mobile crowdsensing is the ability of the network organizer to collect data from the sensors on each user’s mobile phone. This, however, violates the data privacy of mobile users and requires additional on-device power consumption for data transmission [192]. Taking the noise monitoring task as an example, continuous collection of ambient sound violates personal data privacy and reduces devices’ battery life. To encourage as many mobile users as possible to contribute their sensory data with high accuracy and reliability, incentive mechanisms become promising methods to balance the users’ additional cost and the overall performance gain of crowdsensing systems [77, 193].
Based on whether financial payment is applied, the incentive mechanisms can be categorized into monetary and non-monetary [77].
  • Monetary Incentives: Paying for sensory data is the most straightforward incentive. Monetary incentives serve as an important motivator when crowdsensing tasks don’t yield immediate benefits for users. Some research has shown that the increasing amount of payment can help the sensing task be completed faster (see, e.g., [194]). Monetary incentives are further divided into static mechanisms and dynamic mechanisms. In the former mechanisms, the reward quantity is pre-determined based on a specific criterion, and this amount remains consistent throughout the entire experiment (see, e.g., [195]). The latter mechanisms allocate a variable budget for each task, contingent on the real-time conditions of the system (see, e.g., [196]).
  • Non-Monetary Incentives: As its name suggests, non-monetary incentives have no financial payment and call for voluntary participation [197]. This severely restricts its scope of application. In some specific tasks, contributing users can benefit from the overall crowdsensing tasks, making it possible to design appropriate incentive mechanisms, including collective motives, social reward, and intrinsic motives and fun [77]. Collective incentives encourage the achievement of common goals, which is the philosophy of crowdsensing, to work together for a common good. However, once the goal is reached, non-participants can also make a profit. Hence, it’s desirable to add additional incentives to avoid “a free ride”. Social rewards are very useful when collecting sensory data from social networks. The rewards refer to the benefits that can be obtained by social networks (e.g., Facebook and TikTok) including interaction, and self-interests (see, e.g., [198, 199]). In non-monetary incentives, the inherent interest and personal fun, like gaming, are the motivation of the majority of participants [78]. This intrinsic reason can attract the target participants to contribute their own data.
@@ III-D4 Applications
By leveraging the rich data collected via mobile crowdsensing, a wide range of applications are empowered. Several typical applications are listed below.
  • Emergency Management and Prevention Applications like accidents and natural disasters, where urgent decision-making and fast response are required [79].
  • Environmental Monitoring & Waste Management Applications such as air quality [80], noise pollution [200], and waste-recycling operations [201], where the resources, infrastructures, and phenomena, that related to people’s daily life are monitored.
  • E-commerce Applications that collect, process, and deliver prices of goods in the real physical world [81].
  • Indoor Localization Applications that provide localization and navigation services when users are in indoor environments [202].
  • Health Care and Wellbeing Applications that diagnose, treat, and prevent illness, diseases, and injuries using the collected crowdsensing data [203].
  • Intelligent Transportation Systems for efficient transportation and traffic management via monitoring traffic, vehicles, public transport, road conditions, and arrival time [204].
  • Mobile Social Networks that build social relations and meet and share data between people with similar interests using mobile devices [205].
  • Public Safety Applications for evaluating the safety exploiting the urban environments [206].
  • Unmanned Vehicles to enable interaction between mobile users and automobile vehicles [207].
  • Context Awareness that combines mobile activity recognition and data mining to detect context for benefiting data accuracy and battery maintenance [208].
@@ IV-B1 General Integrated Wireless Sensing and Communication
In a DFRC system, radar and communication can be simultaneously performed and jointly designed via sharing hardware platforms and spectrum resources. However, different from communication systems, radar systems have uncooperative targets and the useful information is in the echo signals [82]. This leads to different design principles. To unify the design of radar and communication, researchers have made efforts to investigate the performance metrics of radar from an information-theoretic view. In [221], the capacity of a round-trip radar channel is defined as the number of distinguishable targets. Furthermore, the authors in [222] point out that the maximum achievable information of detected targets of a radar system is the mutual information of the transmit signal and the echo waveform. Thereby, the target of echo signal processing is to reduce the uncertainty of detected targets for enhancing the detection precision. The remaining part of the uncertainty is lower-bounded by the Cramér-Rao bound (CRB) [223]. On this basis, the performance tradeoff between communication and radar is investigated [222, 224, 225].
@@ IV-C Integrated Multi-Modal Sensing and Communication
Although integrated wireless sensing and communication have achieved significant development, such techniques cannot fully utilize the massive non-RF sensors and the corresponding sensory data [38]. What’s more, in future wireless networks, the electromagnetic environment may change rapidly, leading to severe frequency selectivity and time selectivity in wireless channels. The performance gain from integrated wireless sensing and communication, however, comes from the electromagnetic wave propagation domain, which cannot provide either complete or fine-grained characterization of the rapidly-changing environment. This leads to little performance enhancement potential [38]. On the contrary, multi-modal sensing systems obtain information of the environment through multiple domains, including radio wave, visible light or infrared light reflections. With the widely deployed multi-modal sensors, the parameters of the physical environment can be acquired more precisely and efficiently, thereby enhancing communication efficiency. On the other hand, with fast and reliable communication networks, multi-modal sensory data can be shared, exchanged, and fused among different devices, which can further significantly enhance the sensing quality and the performance of downstream tasks such as AI model training and inference. To this end, the research paradigm of integrated multi-modal sensing and communication is proposed [38]. In the sequel, the challenges and techniques are first elaborated. Then, a promising research topic of integrated multi-modal sensing and communication called haptic communications, is introduced [85], which is regarded as a key technology in 6G.
@@ IV-C1 Challenges and Techniques
The efficient implementation of integrated multi-modal sensing and communication faces two main kinds of challenges. On one hand, the sensory data obtained from non-RF sensors are usually in an abstract format. This makes it difficult to be integrated with communication, thereby resulting in a lack of processing methods with interpretability and expansibility. On the other hand, there is currently a lack of datasets that contain aligned multi-modal sensory data and wireless channel information. As a result, the mapping relation between multi-modal sensing and communication remains an unexplored issue.
@@
The techniques of integrated multi-modal sensing and communication can be grouped into three types. The first type targets constructing datasets for building up the relation between multi-modal sensory data and communication environments. The second type aims at exploiting the multi-modal sensory data and machine learning algorithms for CSI estimation. The last type enhances the performance of multi-modal-sensing-based intelligent tasks.
@@
  • Dataset Construction: Since there are various kinds of multi-modal sensors, the obtained multi-modal sensory data are in an abstract format, which results in a lack of processing methods with interpretability and expansibility. Besides, the mapping relation between multi-modal sensing and communication is unexplored. To tackle these issues, a good dataset of integrated multi-modal sensing and communication should satisfy the following requirement [38]. To begin with, the sensory data and wireless channel data should be aligned to achieve mutual promotion between each other. Then, the sensory data should be collected from various devices in order to enlarge the sensing ranges, thereby enhancing the scalability and generality of the supported applications. Moreover, the dataset should be obtained from real scenarios to make sure that its supported research and projects have practical and high-generation ability. To this end, a vision-aided wireless communication dataset is constructed using 3D modeling and ray-tracing simulators [255], where RGB images and depth maps are created from the visual instances, and the wireless features are obtained with sinusoid waveform generator. By exploiting the dataset in [255], a machine-learning-based framework is proposed in [256] for enabling proaction in high-frequency wireless networks. The authors in [257] develop a simulation dataset called M3SC in dynamic vehicular communication networks, where the multi-modal sensory data and wireless channel data are aligned and massive MIMO and millimeter wave (mmWave) communications are considered. For better characterizing the practical imperfections, a large-scale integrated multi-modal sensing and communication dataset is proposed in [258], where the data are collected from real-world scenarios with mmWave receivers, RGB cameras, 3D LiDARs, RF radars and GPS receivers.
@@
  • CSI Estimation for Communication Enhancement: Multi-modal sensory data include information on key elements in the environments such as the position of scatters, that are highly relevant to CSI. Therefore, using multi-modal sensory data for CSI estimation has attracted much focus. However, the mapping relation between multi-modal sensory data and communication characteristics is difficult to obtain. Many works manage to build up such a relation via using deep neural networks (e.g., [84, 259, 260]). Another research direction targets enhancing the communication performance in MIMO systems, where beam should be aligned to increase the received SNR. Multi-modal sensory data has been useful for predictive beamforming, where the position of a target is predicted with the assistance of multi-modal sensors so as to perform the corresponding beamforming in advance (e.g., [261, 262]).

### v2

@@ III-C2 Applications
Multi-modal sensing has empowered various kinds of applications. Via the utilization of audio and visual data, the performance of speech recognition and synthesis [148], human action classification [149], multimedia event detection [150], emotion recognition [151], can be greatly enhanced compared with using unimodal information. By utilizing the modality translation techniques, many media description applications have been well developed, including image and video captioning [152], visual question-answering [153], etc. By creating structured representations of multi-modal data, multimedia content indexing, and retrieval applications are enabled, including automatic shot-boundary detection [154] and video summarization [155]. Furthermore, multi-modal data can enhance the understanding, analysis, and decision-making processes in the domain of healthcare, fitness, and general wellness [156, 157]. Autonomous systems utilize various kinds of sensory data, including radar, stereoscopic visible-light cameras, light detection and ranging (LiDAR), infrared cameras, a global positioning system (GPS), and audio for enhancing the automation capability [158, 159].
@@ III-D2 Key Techniques
  • Data Collection: In mobile crowdsensing systems, data collection requires the engagement of massive user participation, where some non-professional users are included. The target is acquiring reliable and trustworthy data via sensor selection. There are two frameworks. One is the domain-specific framework designed for a dedicated task such as intelligent transportation and health care [34]. Its shortage lies in the co-existence of many tasks or domains degrading the efficiency and reliability of the overall crowdsensing system. The other is a general-purpose data collection framework. It supports many tasks simultaneously by maximizing a set of parameters or minimizing the costs [167].
  • Data Storage: The practical storage location in crowdsensing systems is task-oriented and is classified into centralized storage and distributed storage. The former stores data in a single location, typically a cloud-based database, which is applied when significant processing or data aggregation is needed, e.g., in urban monitoring and price comparison [34]. This approach requires high communication throughput for data transmission and violates the data privacy of users but can enjoy a low processing latency. Distributed storage is employed for delay-tolerant applications such as mapping air quality and fog computing, but has low communication cost and can keep users’ data privacy [168].
  • Signal Processing: Signal processing in crowdsensing refers to the manipulation and analysis of the collected data. It involves three sequential processes, i.e., pre-processing, analytics, and post-processing [34]. The main target of pre-processing is filtering and denoising by removing irrelevant and redundant information, suppressing noise, aggregating data, and compressing data size to save storage volume [169]. Analytics aims to extract meaningful information from the collected data. Machine learning and data mining are common approaches for information inference, pattern recognition, or future trend prediction. For real-time tasks, high computational resources are needed for fast decision-making and response [34]. After analytics, post-processing can be conducted in an offline way for statistical or predictive analysis [170].
@@
    Task assignment involves two aspects. One is task partitioning, which divides complex crowdsensing tasks into smaller and manageable sub-tasks, that can be completed by different devices [171]. For example, in [172], the overall target area is divided into multiple distinct sub-areas, with each being assigned to one device for sensing. The spatial and temporal correlation of different devices’ sensory data are leveraged. The other task assignment technique is task allocation, which involves distributing tasks to users with the goal of reducing resource costs and enhancing task performance. This technique is categorized into two groups, i.e., centralized methods and distributed methods. In the former, a central coordinator assigns the tasks to users or devices, leading to an enhanced systematic performance but requiring global information (e.g., [172]). In the latter, each device makes its own decision during the sensing task, resulting in better data privacy preservation (e.g., [173]).
    Task scheduling in mobile crowdsensing refers to the way that mobile devices collect and contribute their sensory data. In proactive scheduling, mobile devices actively obtain sensory data without a pre-assigned task and can decide where and when to contribute their data. Proactive scheduling is commonly conducted for people-centric tasks (e.g., [174]). In reactive scheduling, the sensing task and target are clearly pre-defined, which is adopted in many works. For example, the authors in [175] propose a minimum energy multi-sensor task scheduling algorithm that coordinates multiple tasks for the same user under the targets of enhancing sensory data quality and minimizing energy consumption.
@@ III-D3 Incentive Mechanisms
The crucial condition for the success of mobile crowdsensing is the ability of the network organizer to collect data from the sensors on each user’s mobile phone. This, however, violates the data privacy of mobile users and requires additional on-device power consumption for data transmission [176]. Taking the noise monitoring task as an example, continuous collection of ambient sound violates personal data privacy and reduces devices’ battery life. On one hand, many techniques have been proposed to enhance data privacy preservation, including data anonymization by removing personally identifiable information before it is shared [177], data encryption to avoid unauthorized access [178], adding noise under the guideline of differential privacy [179], and designing secure data-sharing protocols and decentralized data processing methods [180]. On the other hand, to encourage as many mobile users as possible to contribute their sensory data with high accuracy and reliability, incentive mechanisms become promising methods to balance the users’ additional cost and the overall performance gain of crowdsensing systems [181, 182]. Based on whether financial payment is applied, the incentive mechanisms can be categorized into monetary and non-monetary [181]. In monetary incentives, paying for sensory data is the most straightforward incentive. Some research has shown that the increasing amount of payment can help the sensing task be completed faster (e.g., [183]). They are further divided into static mechanisms and dynamic mechanisms. In the former mechanisms, the reward quantity is pre-determined based on a specific criterion, and this amount remains consistent throughout the entire experiment (e.g., [184]). The latter mechanisms allocate a variable budget for each task, contingent on the real-time conditions of the system (e.g., [185]). Non-monetary incentives have no financial payment and call for voluntary participation [186]. This severely restricts its scope of application. In some specific tasks, contributing users can benefit from the overall crowdsensing tasks, making it possible to design appropriate incentive mechanisms, including collective motives, social reward, and intrinsic motives and fun [181]. Collective incentives encourage the achievement of common goals, which is the philosophy of crowdsensing, to work together for a common good. However, once the goal is reached, non-participants can also make a profit. Hence, it’s desirable to add additional incentives to avoid “a free ride”. In non-monetary incentives, the inherent interest and personal fun, like gaming, are the motivation of the majority of participants [187].
@@ III-D4 Applications
By leveraging the rich data collected via mobile crowdsensing, a wide range of applications are empowered, including emergency management and prevention for accidents and natural disasters [188], environmental monitoring & waste management applications such as air quality [166], noise pollution [189], and waste-recycling operations [190], E-commerce applications for collecting, processing, and delivering prices of goods in the real physical world [191], mobile social networks that build social relations and meet and share data between people with similar interests using mobile devices [192], public safety applications for evaluating the safety exploiting the urban environments [193], health care and wellbeing applications [194], intelligent transportation systems [195], etc.
III-E Discussion
All of the above three ISC technologies have vast applications, especially in the era of AI. Each of them has its own pros and cons. Wireless sensing enjoys the benefits of simple data format, mature signal processing techniques, long-detection distance, high penetration properties, and ease of being equipped in vehicles. The key technical issues of wireless sensing include architecture design, waveform design, and signal processing to overcome the challenges caused by the hostile wireless environments. Multi-modal sensing is more robust and can enhance task performance by exploiting and fusing the information from data in multiple modalities as already mentioned. However, the key issues like representation, alignment, and fusion, call for complicated AI-based algorithms. Besides, there is a lack of theoretical methods to guide its design. Mobile crowdsensing has the advantages of abundant data, high coverage and scalability, low cost, and fast response. However, it takes the risk of violating data privacy and thus requires more efforts to encourage users to contribute data.
@@ IV-B1 General Integrated Wireless Sensing and Communication
In a DFRC system, radar and communication can be simultaneously performed and jointly designed via sharing hardware platforms and spectrum resources. However, different from communication systems, radar systems have uncooperative targets and the useful information is in the echo signals [208]. This leads to different design principles. To unify the design of radar and communication, researchers have made efforts to investigate the performance metrics of radar from an information-theoretic view. In [209], the capacity of a round-trip radar channel is defined as the number of distinguishable targets. Furthermore, the authors in [210] point out that the maximum achievable information of detected targets of a radar system is the mutual information of the transmit signal and the echo waveform. Thereby, the target of echo signal processing is to reduce the uncertainty of detected targets for enhancing the detection precision. The remaining part of the uncertainty is lower-bounded by the CRB [211]. On this basis, the performance tradeoff between communication and radar is investigated [210, 212, 213].
@@ IV-C Integrated Multi-Modal Sensing and Communication
Although integrated wireless sensing and communication have achieved significant development, such techniques cannot fully utilize the massive non-RF sensors and the corresponding sensory data [39]. What’s more, in future wireless networks, the electromagnetic environment may change rapidly, leading to severe frequency selectivity and time selectivity in wireless channels. The performance gain from integrated wireless sensing and communication, however, comes from the electromagnetic wave propagation domain, which cannot provide either complete or fine-grained characterization of the rapidly changing environment. This leads to little performance enhancement potential [39]. On the contrary, multi-modal sensing systems obtain information of the environment through multiple domains, including radio waves, visible light, or infrared light reflections. With the widely deployed multi-modal sensors, the parameters of the physical environment can be acquired more precisely and efficiently, thereby enhancing communication efficiency. On the other hand, with fast and reliable communication networks, multi-modal sensory data can be shared, exchanged, and fused among different devices, which can further significantly enhance the sensing quality and the performance of downstream tasks such as AI model training and inference. To this end, the research paradigm of integrated multi-modal sensing and communication is proposed [39]. In the sequel, the challenges and techniques are first elaborated. Then, a promising research topic of integrated multi-modal sensing and communication called haptic communications, is introduced [244], which is regarded as a key technology in 6G.
@@ IV-C1 Challenges and Techniques
The efficient implementation of integrated multi-modal sensing and communication faces two main kinds of challenges [39]. On one hand, the obtained multi-modal sensory data are in various formats and differ in physical properties like resolution, bandwidth, and latency. To characterize the rapidly changing wireless channel information, it calls for fast and accurate data collection and processing methods. On the other hand, there is currently a lack of datasets that contain aligned multi-modal sensory data and wireless channel, especially for customized scenarios, due to labor and cost concerns. As a result, the mapping relation between multi-modal sensing and communication remains an unexplored issue.
@@
The techniques of integrated multi-modal sensing and communication can be grouped into three types. The first type targets constructing datasets for building up the relation between multi-modal sensory data and communication environments. The second type aims to exploit the multi-modal sensory data and machine learning algorithms for CSI estimation. The last type enhances the performance of multi-modal-sensing-based intelligent tasks.
@@
  • Dataset Construction: As mentioned, the multi-modal sensory data collected from different types sensors have heterogeneous formats and physical properties which results in a lack of processing methods with interpretability and expansibility. Besides, the mapping relation between multi-modal sensing and communication is unexplored. To tackle these issues, a good dataset of integrated multi-modal sensing and communication should satisfy the following requirement [39]. To begin with, the sensory data and wireless channel data should be aligned to achieve mutual promotion between each other. Then, the sensory data should be collected from various devices in order to enlarge the sensing ranges, thereby enhancing the scalability and generality of the supported applications. Moreover, the dataset should be obtained from real scenarios to make sure that its supported research and projects have practical and high-generation ability. To this end, a vision-aided wireless communication dataset is constructed using 3D modeling and ray-tracing simulators [245], where RGB images and depth maps are created from the visual instances, and the wireless features are obtained with sinusoid waveform generator. By exploiting the dataset in [245], a machine-learning-based framework is proposed in [246] for enabling proaction in high-frequency wireless networks. The authors in [247] develop a simulation dataset called M3SC in dynamic vehicular communication networks, where the multi-modal sensory data and wireless channel data are aligned and massive MIMO and millimeter wave (mmWave) communications are considered. For better characterizing the practical imperfections, a large-scale integrated multi-modal sensing and communication dataset is proposed in [248], where the data are collected from real-world scenarios with mmWave receivers, RGB cameras, 3D LiDARs, RF radars and GPS receivers.
@@
  • CSI Estimation for Communication Enhancement: Multi-modal sensory data include information on key elements in the environments such as the position of scatters, that are highly relevant to CSI. Therefore, using multi-modal sensory data for CSI estimation has attracted much focus. However, the mapping relation between multi-modal sensory data and communication characteristics is difficult to obtain. Many works manage to build up such a relation via using DNNs (e.g., [249, 250, 251]). Another research direction targets enhancing the communication performance in MIMO systems, where beam should be aligned to increase the received SNR. Multi-modal sensory data has been useful for predictive beamforming, where the position of a target is predicted with the assistance of multi-modal sensors so as to perform the corresponding beamforming in advance (e.g., [252, 253]). Particularly, the use of multi-modal data for position prediction enjoys the benefit of higher robustness and accuracy due to the common and supplementary information provided by data in different modalities.